OpenStudy (anonymous):

The organizers of a television show must select 5 people to participate in the show. The participants will be selected from a list of 30 people who have written in to the show. If the participants are selected randomly, what is the probability that the 5 youngest people will be selected?

OpenStudy (kropot72):

The probability of selecting the first of the five youngest people is 5/30. The probability of selecting the second of the five youngest people is 4/29. Using this reasoning regarding the selection of the remaining three of the five youngest people, the required probability is given by: \[\frac{5}{30} \times \frac{4}{29} \times\frac{3}{28}\times\frac{2}{27}\times\frac{1}{26}=you\ can\ calculate\]